# Quantum nonlinear effects in the number-conserving analogue gravity of Bose-Einstein condensates

### v1

@@ VI Conclusion
The intrinsic nonlinearity of every interacting quantum field theory (classically and in the hydrodynamical limit a fluid-dynamical system with a given equation of state) requires in general that one goes beyond the standard linearized Unruh paradigm when formulating an analogue gravity scenario. To account self-consistently for effects which stem from the nonlinearity, to second (or higher) order in the fluctuations, one needs to take backreaction of the fluctuations onto the condensate into account. We have demonstrated, to second order in a number-conserving approach to the quantum many-body problem in powers of $N^{1-/2}$, that this engenders, when enforcing Lorentz invariance of the wave equation for the second-order fluctuations, corrections to the Unruh paradigm metric (which is by definition seen by the first-order fluctuations). To higher order in $N^{-1/2}$, quantum nonlinearity, in addition, creates emergent vector fields. We have also demonstrated that number conservation implies the presence of dissipation for excitations above background-II.
While probing these correcting nonlinear effects on the standard analogue gravity Unruh paradigm will be experimentally challenging, they might in future experiments become testable by increasing, in a highly controlled way, the contact coupling $U$ via Feshbach resonances. This, then, takes the detail to which quantum field dynamics in curved spacetime, and thus the background field concept itself, can be tested to another level of accuracy. In experimental terms, this is contingent on the ability to measure higher-order correlation functions than those of the (standard) density-density variety, as experimentally demonstrated, e.g., in Ref. [50].
@@ VII ACKNOWLEDGMENTS
We thank Satadal Datta and Caio César Holanda Ribeiro for helpful discussions. K. P. would like to thank Arun and Kuntal Pal for important discussions. This work has been supported by the National Research Foundation of Korea under Grants No. 2017R1A2A2A05001422 and No. 2020R1A2C2008103.
@@
We show here how the quantum backreaction term $\phi_{2}$ affects the evolution of a quantum-backreacted condensate in a box trap. According to the fluid dynamical description of the BEC, which we have employed here, the primary two equations governing the dynamics are (19) and (20). These two equations in terms of the classical fluid variables are different from those obtained from the solutions of the GPE (5), due to the presence of $f_{1}$ and $f_{3}$, which encapsulate the contribution from the depletion cloud. Here, our primary goal is an estimation of the size and development of these contributions in a realistic model of BEC which accounts for backreaction. We will to this end consider the model studied previusly in [35], where an initially noninteracting Bose gas of uniform density at rest, trapped in a hard-walled container (one-dimensional box trap) of total length $\ell$ extending from $x=-\ell/2$ to $x=+\ell/2$, is subjected to an interaction quench at $t=0$. For $t<0$, the gas is described as solution of the GPE $\phi_{0}$ such that $\rho_{0}=|\phi_{0}|^{2}=$ constant and is at rest. In the noninteracting regime ($U=0$), the density of the depleted cloud $\rho_{\hat{\chi}}=0$, and consequently backreaction is also not present, i.e., $\phi_{2}=0$. For $t\geq 0$, according to the backreaction scheme developed in [35], the leading-order condensate order parameter $\phi_{0}$ remains unchanged, and quantization of the field $\hat{\chi}$, along with the condensate-correction described by $\phi_{2}$, describes the coupled evolution of the condensate and non-condensate parts.
In general, the coupled dynamics of $\phi_{c}$ and $\hat{\chi}$ are difficult to solve for analytically; however, due to the assumption of a constant order parameter to leading order in $\mathcal{O}(N^{1-/2})$, it is possible to exactly quantize and obtain $\hat{\chi}$ for the model considered, and similarly, a solution for $\phi_{2}$ can be obtained fully analytically [37, 36]. To compute the function $f_{1}(x,t)$ related to the real part of the anomalous correlator, we resort to the continuity equation satisfied by the background $\phi_{c}$, Eq. (16), to obtain
@@
Since the analytical solution for $\phi_{c}$ in terms of $\phi_{0}$ and $\phi_{2}$ is known, it is possible to compute $f_{1}(x,t)$ from (58), taking into account that $\phi_{0}=\exp(-i\mu t)\sqrt{\rho_{0}}$, where $\mu$ is the chemical potential, and $\rho_{0}$ is constant. For being self-contained, we quote here in full the form of density $\rho_{\phi_{2}}$, and current $J_{\phi_{2}}$ for $t\geq 0$ from [35]:
$\rho_{\phi_{2}}$ $=$ $2\mbox{Re}[\exp(i\mu t)\sqrt{\rho_{0}}\phi_{2}]$ (59)
$=$ $-\frac{t^{2}}{\ell}-\frac{1}{4\ell}\sum_{n=1}^{\infty}\frac{(-1)^{n}}{\omega_{n}^{2}}\left\{2(-1)^{n}[1-\cos(2\omega_{n}t)]+\cos(2k_{n}x)\left[\frac{2-k_{n}^{2}}{1+k_{n}^{2}}+2\cos(2\omega_{n}t)-\frac{k_{n}^{2}+4}{k_{n}^{2}+1}\cos(\omega_{2n}t)\right]\right\}\!,$
$J_{\phi_{2}}$ $=$ $\mbox{Im}\ [\phi_{0}^{*}\partial_{x}\phi_{2}]=\frac{-2}{\ell}\sum_{n=1}^{\infty}\frac{(-1)^{n}\sin(2k_{n}x)}{k_{n}}\left[\frac{\sin(2\omega_{n}t)}{2\omega_{n}}-\frac{\sin(\omega_{2n}t)}{\omega_{2n}}\right].$ (60)
@@
These expressions can be obtained from the backreaction field $\phi_{2}$, which was solved for exactly in [35]. Here, we have indicated mode frequencies $\omega_{n}=\sqrt{k_{n}^{2}(k_{n}^{2}/4+1)}$, and wavevectors $k_{n}=n\pi/\ell$, which are a result of imposing the Neumann boundary conditions at $x=\pm\ell/2$.
Figure A.1: (a) Temporal evolution of the spatial dependence of $f_{1}$ for a condensate of total length $\ell=40$, extending from $x=-20$ to $x=20$. Red: $t=0.5$, Black: $t=1$, Blue: $t=2$, Green: $t=3$, Magenta: $t=4$. (b) Evolution of $f_{1}$ for a condensate of larger size $\ell=100$; instants of time indicated with identical colors as in (a). We have verified that the plots converge for mode numbers $n\gtrsim 300$.
@@
We present in Figs. A.1 selected plots of the variation of $f_{1}(x,t)$ with $x$ at various instants $t\geq 0$, where $t$ is in units of the chemical potential of the background with respect to the lowest order density, $U\rho_{0}$. We note that for all profiles the function $f_{1}<0$, which confirms that there is an out-going particle flux leaving the condensate (particle loss by the condensate). After the system is quenched from a noninteracting to an interacting regime, $f_{1}$ becomes nonzero, following a pattern similar to that of $\rho_{\hat{\chi}}$ [35], and such that the growth in the bulk of the system is not sensitive to the boundary region.
Figure A.2: Temporal evolution of $f_{1}$ for a condensate of total length $\ell=40$ at several locations $x$, parameters as in Fig. A.1 (a). Red: $x=5$, Black: $x=10$, Blue: $x=15$, Green: $x=17$, Magenta: $x=19$. Note that just after the interaction is turned on, there is a universal growth of particle loss from the condensate, irrespective of the position in the trap.
@@
We also plot in Fig. A.2, the evolution of $f_{1}$ with time for different locations inside the trap. We infer that the growth of $f_{1}$ at different $x$, just after the interaction is turned on, follows the same universal pattern. It is dominated by a quadratic growth, which is the contribution of the zero mode contained in (59). This growth is universal irrespective of the location inside the trap. Furthermore, the evolution is drastically different depending on the location in the trap. Near the center of the gas, it is oscillatory. However, close to the hard wall, the contribution to the backreaction quickly saturates with time. These observations are consistent with the observations made in Ref. [36], that as the interaction is suddenly turned on, the energy exchange of the condensate with the quasiparticle excitation cloud initially occurs very fast, and then saturates to an oscillatory behavior, within a time scale depending on the position in the trap.

### v2

@@ VI Conclusion
The intrinsic nonlinearity of every interacting quantum field theory (classically and in the hydrodynamical limit a fluid-dynamical system with a given equation of state) requires in general that one goes beyond the standard linearized Unruh paradigm when formulating an analogue gravity scenario. Specifically, to account self-consistently for effects which stem from the nonlinearity, to second (or higher) order in the fluctuation amplitude, one needs to take backreaction of the fluctuations onto the condensate into account. We have demonstrated in our approach to the quantum many-body problem that this engenders, when enforcing Lorentz invariance of the wave equation for the second-order fluctuations, to leading order in the many-body expansion in powers of $N^{1-/2}$, corrections to the Unruh paradigm metric (which is by definition seen by the first-order fluctuations). To higher order in $N^{-1/2}$, quantum nonlinearity, in addition, creates emergent vector fields. We have furthermore shown that number conservation implies the potential presence of a imaginary part of the frequency for excitations above the background-II for second-order perturbations.
While probing these correcting nonlinear effects on the standard analogue gravity Unruh paradigm will be experimentally challenging, they might in future experiments become testable. For example by increasing, in the highly controlled way feasible in the quantum optical context, the contact coupling $U$ via Feshbach resonances to simulate quantum many-body effects on a FLRW background as discussed above. This, then, takes the detail to which quantum field dynamics in curved spacetime, and thus the background field concept itself, can be tested to another level of accuracy. In experimental terms, this is contingent on the ability to measure higher-order correlation functions than those of the (standard) density-density variety, as experimentally demonstrated, e.g., in Ref. [53].
@@ VII ACKNOWLEDGMENTS
We thank Satadal Datta and Caio César Holanda Ribeiro for helpful discussions. KP would like to acknowledge Arun Rana and Kuntal Pal for instructive conversations. This work has been supported by the National Research Foundation of Korea under Grants No. 2017R1A2A2A05001422 and No. 2020R1A2C2008103.
@@
We show here how the quantum backreaction term $\phi_{2}$ affects the evolution of a quantum-backreacted condensate. According to the fluid dynamical description of the BEC, which we have employed here, the primary two equations governing the dynamics are (21) and (22). These two equations in terms of the classical fluid variables are different from the continuity and Euler equations obtained from the GPE (5) due to the presence of $f_{1}$ and $f_{3}$, which encapsulate the backreaction contribution from the depletion cloud. Here, our primary goal is an estimation of the size and development of these contributions in a realistic model of a BEC which accounts for backreaction. We will to this end consider the model studied previously in [36], where an initially noninteracting Bose gas of uniform density at rest, trapped in a hard-walled one-dimensional container (i.e., in a one-dimensional box trap with no lateral movement permitted by the trapping geometry) of total length $\ell$ extending from $x=-\ell/2$ to $x=+\ell/2$, is subjected to an interaction quench at $t=0$. For $t<0$, the gas is described as solution of the GPE $\phi_{0}$ such that $\rho_{0}=|\phi_{0}|^{2}=$ constant and is at rest. In the noninteracting regime ($U=0$), the density of the depleted cloud $\rho_{\hat{\chi}}=0$, and consequently backreaction is also not present, i.e., $\phi_{2}=0$. For $t\geq 0$, the leading-order condensate order parameter $\phi_{0}$ remains unchanged, and quantization of the field $\hat{\chi}$, along with the condensate-correction described by $\phi_{2}$, describes the coupled evolution of the condensate and non-condensate parts fully analytically [38, 37].
To compute the function $f_{1}(x,t)$ related to the real part of the anomalous correlator, we resort to the continuity equation satisfied by the background $\phi_{c}$, Eq. (18), to obtain
@@
We then use that for the solution of Ref. [38] we have $\rho_{\phi_{c}}=\rho_{\phi_{0}}+\rho_{\phi_{2}}+\mathcal{O}(N^{-1/2})$ and $J_{\rm c}=J_{\phi_{0}}+J_{\phi_{2}}+\mathcal{O}(N^{-1/2})$, where $J_{\phi_{0}}=\mbox{Im}[\phi^{*}_{0}\partial_{x}\phi_{0}]$ and
$\rho_{\phi_{2}}$ $=2\mbox{Re}\ [\phi_{0}^{*}\phi_{2}],$ (61)
$J_{\phi_{2}}$ $=\mbox{Im}\ [\phi_{0}^{*}\partial_{x}\phi_{2}+(\partial_{x}\phi_{0})\phi_{2}^{*}].$ (62)
Since the analytical solution for $\phi_{c}$ in terms of $\phi_{0}$ and $\phi_{2}$ is known, it is possible to compute $f_{1}(x,t)$ from (60), setting $\phi_{0}=\exp(-i\mu_{0}t)\sqrt{\rho_{0}}$, where $\mu_{0}=U\rho_{0}$ is the (lowest order) chemical potential, and the lowest-order density $\rho_{0}$ is constant, which defines the bulk coherence length scale via $1/\xi_{0}^{2}=U\rho_{0}=c_{s0}^{2}$, so that all quantities are expressed in powers of $\xi_{0}$.
For being self-contained, we quote here in full the form of the second-order contributions to density $\rho_{\phi_{2}}$ and current $J_{\phi_{2}}$, for $t\geq 0$ (where $t$ is in units of $\xi_{0}^{2}$), from Ref. [36]
$\rho_{\phi_{2}}$ $=$ $-\frac{t^{2}}{\ell}-\frac{1}{4\ell}\sum_{n=1}^{\infty}\frac{(-1)^{n}}{\omega_{n}^{2}}\left\{2(-1)^{n}[1-\cos(2\omega_{n}t)]+\cos(2k_{n}x)\left[\frac{2-k_{n}^{2}}{1+k_{n}^{2}}+2\cos(2\omega_{n}t)-\frac{k_{n}^{2}+4}{k_{n}^{2}+1}\cos(\omega_{2n}t)\right]\right\}\!,$ (63)
$J_{\phi_{2}}$ $=$ $\frac{-2}{\ell}\sum_{n=1}^{\infty}\frac{(-1)^{n}\sin(2k_{n}x)}{k_{n}}\left[\frac{\sin(2\omega_{n}t)}{2\omega_{n}}-\frac{\sin(\omega_{2n}t)}{\omega_{2n}}\right].$ (64)
@@
These expressions can be obtained from the backreaction field $\phi_{2}$, which was solved for exactly in [36]. Here, we have indicated mode frequencies $\omega_{n}=\sqrt{k_{n}^{2}(k_{n}^{2}/4+1)}$, and wavevectors $k_{n}=n\pi/\ell$, which are a result of imposing Neumann boundary conditions at $x=\pm\ell/2$.
@@
We present in Figs. A.1 (a) and (b) selected plots of the variation of $f_{1}(x,t)$ with $x$ at various instants $t\geq 0$. We note that for all profiles the function $f_{1}<0$, which confirms that there is an outgoing particle flux leaving the condensate (particle loss by the condensate). After the system is quenched from a noninteracting to an interacting regime, $f_{1}$ becomes nonzero, following a pattern similar to that of $\rho_{\hat{\chi}}$ [36], and such that the growth in the bulk of the system is not sensitive to the boundary region.
@@
We also plot in Fig. A.2 the evolution of $f_{1}$ with time for different locations inside the trap. We infer that the growth of $f_{1}$ at different $x$, just after the interaction is turned on, follows the same universal pattern. It is dominated by a quadratic growth, which is the contribution of the zero mode contained in (63). This growth is universal irrespective of the location inside the trap. Furthermore, the evolution is drastically different depending on the location in the trap. Near the center of the gas, it is oscillatory. However, close to the hard wall, the contribution to the backreaction quickly saturates with time. These observations are consistent with the observations made in Ref. [37], that as the interaction is suddenly turned on, the energy exchange of the condensate with the quasiparticle excitation cloud initially occurs very fast, and then saturates to an oscillatory behavior, within a time scale depending on the position in the trap.
Figure A.1: (a) Temporal evolution of the spatial dependence of $f_{1}$ for a condensate of total length $\ell=40$, extending from $x=-20$ to $x=20$. Red: $t=0.5$, Black: $t=1$, Blue: $t=2$, Green: $t=3$, Magenta: $t=4$. (b) Evolution of $f_{1}$ for a condensate of larger size $\ell=100$; instants of time indicated with identical colors as in (a). We have verified that the plots do not visibly differ for mode numbers $n\gtrsim 300$; $x$ is in units of $\xi_{0}$ and $f_{1}$ is in units of $1/\xi_{0}^{3}$.
Figure A.2: Temporal evolution of $f_{1}$ for a condensate of total length $\ell=40$ at several locations $x$, parameters as in Fig. A.1 (a). Red: $x=5$, Black: $x=10$, Blue: $x=15$, Green: $x=17$, Magenta: $x=19$. Note that just after the interaction is turned on, there is a universal growth of particle loss from the condensate, irrespective of the position in the trap; $x$ is in units of $\xi_{0}$ and $f_{1}$ is in units of $1/\xi_{0}^{3}$.